Directions are to Solve and I don't even know where to start. I thought to set the equation equal to zero, but very confused, can someone please help me??
w^2 - 7w - 18 = 0

Answer by scott8148(6628)

factoring ___ (w - 9) (w + 2) = 0

w - 9 = 0 ___ w = 9

w + 2 = 0 ___ w = -2

two solutions for w